OpenStudy (anonymous):

A glass tube of cross sectional area 10^-4 m^2 is partially filled with water. An oil with a density of 800kg/m3 is poured in the tube and floats on top of the water. The height of the oil above the water surface is 8cm. a)What is the change in pressure at a depth of 10cm below the water surface? Answer says 0.6 kilopascals but I don't know why. b) Would the pressure at this depth increase or decrease if you add more oil to the tube and why?

OpenStudy (anonymous):

Well before it was just the atmospheric pressure (if it's open) plus the water. Now it's the oil + atmospheric pressure + the water already in the glass. Is oil denser than air? No. So the pressure must increase.

OpenStudy (anonymous):

\[P(10 cm) = P(atm) + P(oil) + P(water)\]

OpenStudy (anonymous):

In which the oil FLOATS 8 cm above the water. It's asking you for the change in pressure at 10 cm. So the change in pressure will be New pressure - old pressure. Therefore, old pressure = rho(water)*g*h(water), new pressure = rho(oil)*g*h(oil) + rho(water)*g*h(oil)

OpenStudy (anonymous):

The height of the water isn't given though..

OpenStudy (anonymous):